OpenStudy (anonymous):

f(x)=x-3 find (f*f)(x)and its domain

OpenStudy (anonymous):

is this times or composition?

OpenStudy (anonymous):

times

OpenStudy (anonymous):

oh. i that case \[f(x)\times f(x) = (x-3)(x-3)=x^2-6x+9\] polynomial so unless otherwise stated domain is all real numbers.